OpenStudy (anonymous):

Imagine you are teaching a fellow student how to solve: 4 + 2(x - 1) = 5x - (x + 10) In your own words, explain the process for solving this equation. Please include your work at each step of the process along with the final answer

OpenStudy (ajprincess):

Remove the brackets both sides. That will give you 4+(2*x)-(2*1)=5x-(1*x)-(1*10) 4+2x-2=5x-x-10 When you simplify 2+2x=4x-10 Nw bring all terms containing x to one side. keeping the numbers in the other side. 2+10=4x-2x 12=2x Now divide both sides by 2 12/2=(2x)/2 6=x Is that clear?